Find the perimeter and area of a right triangle if the shortest side is 12 in. and the longest side is 37 in. Include correct units with each part of your solution.

Found 2 solutions by mananth, dabanfield:
Answer by mananth(16949)

since it is right triangle apply pythagoras rheorem
hyp^2-side1^2 = side2^2
37^2-12^=side2^2
sqrt(37^2-12^2) = side2
35 = side2
side1 = 12 inches
side2 = 35 inches
hypotenuse = 37 inches
perimeter = 84 inches
area of right triangle = 1/2 * side1 * side2
A= 1/2 *12*35
A= 210 square inches





Answer by dabanfield(803)

Find the perimeter and area of a right triangle if the shortest side is 12 in. and the longest side is 37 in. Include correct units with each part of your solution.

Let x be the third side. Then by the Pythagorean Theorem we have:
x^2 + 12^2 = 37^2
x^2 + 144 = 1369
x^2 = 1369-144
x^2 = 1225
x = 35
The perimeter Of the triangle is 12+35+37 = 84
The area, using the formula Area = (base*height)/2, is:
(35*12)/2 = 210